# R[$\arraybackslashb#1 \newcolumntypeL$$\arraybackslashb#1 \newcolumntypeC$$\arraybackslashb#1$

### v1

@@ 1. Introduction
During the recent years many researchers are interested in spectral graph theory [AtDa, AtEnGo, AtEnGoJa1, AtEnGoJa2, Ay, Co1, Co2, Li, Sh, BoGo, GoTr, Go2, EnJa] mainly the matter of essential self-adjointess of the discrete Laplacian; see [Mi2, Mi1, MiFr, Sc, Go1, CoTo1, CoTo2].
Spectral graph theory has been extended beyond one-dimensional structures to higher-order discrete objects such as simplicial complexes. While graphs correspond to 1-dimensional simplicial complexes, triangulations and higher-dimensional networks naturally lead to $n$-simplicial complexes, where topological and geometric features interact in nontrivial ways, see [AnAychTo, DaJo, DuKlMa, Ch1, AnTo, AtBaDaEn, EnJa].
@@
In the graph case ($n=1$), [AnTo] introduced the notion of $\chi$-completeness, a geometric criterion ensuring ESA of the Gauss–Bonnet operator $D=d+\delta$ and the associated Laplacian $L=D^{2}$. This idea was later extended to 2-dimensional triangulations [BaBeJe, Ch1] and magnetic graphs [AtBaDaEn].
@@
These findings highlight a fundamental principle: essential self-adjointness in simplicial complexes must be analyzed both spatially and dimensionally. A defect in a single degree or region can propagate and obstruct the global spectral theory, unless compensated by analytic or geometric control. The paper is structured as follows. Sections 2 and 3 introduce the basic definitions of weighted simplicial complexes, function spaces, and the Gauss–Bonnet operator. Section 4 defines the three notions of $\chi$-completeness, emphasizing the role of plateau cut-off functions and states and proves the four main theorems. Section 5 provides concrete examples and counterexamples and discusses the implications and outlines future directions. Section 6 gives a comparison with [BaGoJe] and proves ESA in the case of a combinatorial $n$-simplicial complex.
@@ 6.2. Weighted $\chi$-Completeness and Spectral Regularization
Finally, since $L=D^{2}$ and $D$ is essentially self-adjoint, it follows that $L$ is also essentially self-adjoint on $\bigoplus_{i=0}^{n}C_{c}^{i}(V)$.
This completes the proof. ∎

### v2

@@ 1. Introduction
During the recent years many researchers are interested in spectral graph theory [AtDa, AtEnGo, AtEnGoJa1, AtEnGoJa2, Ay, Co1, Co2, Li, Sh, BoGo, GoTr, Go2, EnJa, BaKe] mainly the matter of essential self-adjointess of the discrete Laplacian; see [Mi2, Mi1, MiFr, Sc, Go1, CoTo1, CoTo2].
Spectral graph theory has been extended beyond one-dimensional structures to higher-order discrete objects such as simplicial complexes. While graphs correspond to 1-dimensional simplicial complexes, triangulations and higher-dimensional networks naturally lead to $n$-simplicial complexes, where topological and geometric features interact in nontrivial ways, see [AnAychTo, DaJo, DuKlMa, Ch1, AnTo, AtBaDaEn, EnJa, BaKe].
@@
In the graph case ($n=1$), [AnTo] introduced the notion of $\chi$-completeness, a geometric criterion ensuring ESA of the Gauss–Bonnet operator $D=d+\delta$ and the associated Laplacian $L=D^{2}$. This idea was later extended to 2-dimensional triangulations [BaBeJe, Ch1] and magnetic graphs [AtBaDaEn]. We also note that a more general framework for essential self-adjointness of Hodge Laplacians on abstract simplicial complexes has been developed in [BaKe], where the authors introduce intrinsic-metric criteria. In particular, Proposition 4.12 and Theorem 4.11 of that work establish essential self-adjointness under the assumption that there exists an intrinsic metric with finite distance balls and bounded jump size. As indicated in the remark following Theorem 4.11 therein and by analogy with Theorem 11.31 in [KLW] this metric condition is expected to be equivalent to $\chi$-completeness in the setting of weighted simplicial complexes. Consequently, our Theorems 4.5 and 4.6 can be viewed as concrete realizations of the more abstract criterion in [BaKe, Prop. 4.12].
However, our work goes beyond this general framework in several directions. First, we introduce and distinguish three distinct notions of $\chi$-completeness global, local by level, and local by region and analyze their precise spectral implications, showing that they form a strict hierarchy with no converse implications. Second, we prove essential self-adjointness under localized defects via a compact coupling argument (Theorem 4.7), a technique absent from the intrinsic-metric approach. Third, we establish a divergence criterion (Theorem 4.10) that guarantees essential self-adjointness even in the absence of $\chi$-completeness, thereby resolving in higher dimensions the open question raised in [BaGoJe] about the necessity of geometric completeness. Finally, our setting is tailored to clique complexes arising from weighted graphs, where cut-off functions are defined on vertices and extended by averaging a natural framework for applications in topological data analysis which allows for explicit energy estimates and concrete counterexamples (Section 5) that illustrate the sharpness of our results.
@@
These findings highlight a fundamental principle: essential self-adjointness in simplicial complexes must be analyzed both spatially and dimensionally. A defect in a single degree or region can propagate and obstruct the global spectral theory, unless compensated by analytic or geometric control.
The paper is structured as follows. Sections 2 and 3 introduce the basic definitions of weighted simplicial complexes, function spaces, and the Gauss–Bonnet operator. Section 4 defines the three notions of $\chi$-completeness, emphasizing the role of plateau cut-off functions and states and proves the four main theorems. Section 5 provides concrete examples and counterexamples and discusses the implications and outlines future directions. Section 6 gives a comparison with [BaGoJe] and proves ESA in the case of a combinatorial $n$-simplicial complex.
@@ 6.2. Weighted $\chi$-Completeness and Spectral Regularization
Finally, since $L=D^{2}$ and $D$ is essentially self-adjoint, it follows that $L$ is also essentially self-adjoint on $\bigoplus_{i=0}^{n}C_{c}^{i}(V)$. This completes the proof. ∎
@@
Acknowledgements: The authors are grateful to Prof. P. Bartmann and Prof. M. Keller for their valuable comments and constructive suggestions, which have contributed significantly to improving this paper.